At 10:30am two planes leave Houston, one flying east at 560km/h and the other flying west at 640km/h. At what time will they be 2100km apart?
Answer by mananth(16949)

one plane speed = 560 km/h
other plane speed = 640km/h
..
they are travelling in opposite directions
so effective speed = 560+640= 1200 km/h
..
distance = 2100
speed = 1200 km/h
time = distance /speed
time 2100/1200
= 1.75 hours = 1 hour 45 minutes
they left at 10.30 am.
add 1 hour 45 minutes
12.15 pm. the time they will be 2100 km apart
CHECK
(7/4 * 640)+(7/4 * 560) = 2100 km